Income statement: Sales                   $10,000 Cost of goods sold        9,200 EBIT     &n

Income statement:

Sales                   $10,000

Cost of goods sold        9,200

EBIT                    $   800

Interest (10%)              400

EBT                     $   400

Taxes (40%)                 160

Net Income              $   240

Balance sheet:

Cash                    $    20

A/R                       1,000

Inventories               5,000

Total C.A.              $ 6,020           Debt          $ 4,000

Net F.A.                  2,980           Equity          5,000

Total Assets            $ 9,000           Total claims  $ 9,000

What is Inventory Turn Over Ratio?

Was the company profitable? Yes/No

What is Return on Equity ?

What is Return on Total Assets?
Answer #1

HI

Here Inventory Turn over ratio = Cost of goods sold/average inventory

=9200/500 = 1.84 times

Since Net Income of company is positive hence company is profitable

Return on equity = Net Income / total equity = 240/5000 = 4.80%

Return on total Asset = Net Income/ Total asset = 240/9000 = 2.67%

Thanks